# On computability of equilibrium states

### v2

@@ Abstract.
Equilibrium states are natural dynamical analogs of Gibbs states in thermodynamic formalism. This paper is devoted to the study of their computability in the sense of Computable Analysis. We show that the unique equilibrium state associated to a pair of a computable, topologically exact, distance-expanding, open map $T\colon X\rightarrow X$ and a computable Hölder continuous potential $\varphi\colon X\rightarrow\mathbb{R}$ is always computable. Furthermore, the Hausdorff dimension of the Julia set and the equilibrium state for the geometric potential of a computable hyperbolic rational map are computable. On the other hand, we introduce a mechanism to establish non-uniqueness of equilibrium states. We also present some computable dynamical systems whose equilibrium states are all non-computable.
@@ 1. Introduction
The decidability problem was the third question of Hilbert’s program proposed by Hilbert in Paris in 1900. In the 1930s, the Halting Problem given by Turing provided a counterexample to this question. Another significant contribution of Turing was the construction of a definite method of computations called Turing machine. Turing machines give a standard model of computation in discrete settings and have become the foundation of modern digital computers. Nowadays, however, most digital computers are used for computations in continuous settings. The research on computations with real numbers began with Turing’s work on the original definition of computable real numbers in 1937 ([Tu37]). The work of Banach and Mazur in 1937 ([BM37]) provided definitions of computability for real objects (such as subsets of $\mathbb{R}^{n}$ and functions $f\colon\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$). The “bit model” has been developed, serving as a foundation of the tradition of Computable Analysis ([Gr55, La55, Ko91, Weih00]). Computable Analysis on dynamical systems. In recent decades, there has been dramatic growth in research on the computability and computational complexity of many objects generated by dynamical systems (see for example, [Ya21] and references therein). Initial value sensitivity and typical instability of many interesting systems imply that what is being observed on the computer screen could be completely unrelated to what was meant to be simulated. This paper is devoted to the study of computability questions for a family of invariant measures called equilibrium states (see Definition 2.4). It expands and continues the line of inquiry initiated by [BBRY11] of the first-named author and Braverman, Rojas, and Yampolsky. Recall that for a rational map $f(z)=P(z)/{Q(z)}$ on the Riemann sphere, where $P(z)$ and $Q(z)$ are mutually prime polynomials, with degree $\deg(f)=\max\{\deg(P),\,\deg(Q)\}\geqslant 2$, the Julia set $\mathcal{J}_{f}$ is defined as the locus of the chaotic dynamics of $f$, namely, the complement of the set where the dynamics of $f$ is stable. Moreover, the $*{}^{*}$-limit of the average of Dirac measures supported on the preimage points of a repelling periodic point $w$ is an $f$-invariant measure $\lambda$ (independent of the choice of $w$) called the Brolin–Lyubich measure. In [BBRY11], a uniform machine was designed to compute the Brolin–Lyubich measure for any given rational map. Together with the existence of polynomials whose coefficients are all computable, but whose Julia sets are non-computable from the groundbreaking works of Braverman and Yampolsky ([BY06, BY09]) dating back to a question of Milnor [BY06, Section 1], the computability of Brolin–Lyubich measures leads to a conflict: heuristically speaking, a measure contains more information than its support, but in Computable Analysis, there exists a computable invariant probability measure whose support is, however, non-computable. Such a conflict can be reconciled by considering these two results as the computability properties of the same physical objects from the geometric perspective and the statistical perspective, respectively. In this era of artificial intelligence and data science, the latter undoubtedly deserves closer investigations. Thermodynamic formalism is a powerful method for creating invariant measures with prescribed local behavior under iterations of the dynamical system. This theory, inspired by statistical mechanics, was created by Ruelle, Sinai, and others in the seventies ([Dob68, Si72, Bo75, Wa82]). Since then, thermodynamic formalism has been applied in many classical contexts (see for example, [Bo75, Ru89, Pr90, KH95, Zi96, MauU03, BS03, Ol03, Yu03, PU10, MayU10]) and has remained at the frontier of researches in dynamical systems. Among many other applications, thermodynamic formalism has, since its very early days, played a central role in the study of statistical properties of dynamical systems. The key objects of investigation in thermodynamic formalism are invariant measures called equilibrium states. More precisely, for a continuous map $T\colon X\rightarrow X$ of a compact metric space $(X,\rho)$, a $T$-invariant Borel probability measure $\mu$ on $X$, and a real-valued continuous function $\varphi\colon X\rightarrow\mathbb{R}$, called potential, one can define the measure-theoretic pressure and the topological pressure to describe the degree of complexity of the given dynamical system from different perspectives. The Variational Principle implies that the supremum over all $T$-invariant Borel probability measures on $X$ of the measure-theoretic pressure is equal to the topological pressure. We say that a Borel probability measure $\mu$ is an equilibrium state of $T$ for $\varphi$ if $\mu$ maximizes the measure-theoretic pressure. In particular, if $\varphi$ is a constant function, then an equilibrium state reduces to a measure of maximal entropy (see Subsection 2.2). In particular, the Brolin–Lyubich measure of a rational map $f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ is the measure of maximal entropy of $f$. Statement of main results. The main theorems of this paper include two positive results on the computability of equilibrium states in uniformly expanding systems, an application to complex dynamics, an application to the classical theory of thermodynamic formalism (independent of the interest from Computable Analysis) on non-uniqueness of equilibrium states, and some counterexamples. Computability of equilibrium states. Our first result concerns the computability of equilibrium states. Here is a list of assumptions used in this context.
@@
    $\varphi\colon X\rightarrow\mathbb{R}$ is a real-valued Hölder continuous potential with some exponent $v_{0}$ and constant $a_{0}$ (in the sense of Definition 2.6).
@@
    $T\colon X\rightarrow X$ is an open, topologically exact, continuous map that is distance-expanding (in the sense of Definition 2.1) with respect to the metric $\rho$ with some constants $\eta$, $\lambda$, and $\xi$.
@@
    $\varphi\colon X\rightarrow\mathbb{R}$ and $T\colon X\rightarrow X$ are both computable (in the sense of Definition 3.9).
By classical results (see for example, Proposition 2.5 and [PU10, Corollary 5.2.14]), the Assumption A (i) through (iii) imply that there exists a unique equilibrium state of $T$ for $\varphi$. The Assumption A (iv) allows us to input the information of the system in the computer. For systems satisfying the above assumptions, we prove the following theorem.
@@
Let the quintet $(X,\,\rho,\,\mathcal{S},\,\varphi,\,T)$ satisfy the Assumption A in Section 1, and $\mu$ be the unique equilibrium state of the map $T$ for the potential $\varphi$. Then $\mu$ is computable.
For a more precise version of Theorem A, see Theorem 1.1 below. The Assumption A requires that the behavior of the dynamical system has nice properties in the whole space. However, generally speaking, the dynamical behavior may be merely good enough in an invariant subspace rather than in the whole space. Hence, we prove Theorem B, a “subspace” version of Theorem A. We give a list of assumptions that describes the systems considered in Theorem B as follows.
@@
    $(D,\,\rho,\,\mathcal{S}^{\prime})$ is a computable metric space, where $\mathcal{S}^{\prime}$ is a uniformly computable sequence of points in the computable metric space $(X,\,\rho,\,\mathcal{S})$. $D$ is a recursively compact set in the computable metric space $(D,\,\rho,\,\mathcal{S}^{\prime})$.
@@
    $\varphi\colon D\rightarrow\mathbb{R}$ is a real-valued Hölder continuous potential with some exponent $v_{0}$ and constant $a_{0}$ (in the sense of Definition 2.6).
@@
    $T\colon X\rightarrow X$ is a map such that $D$ is completely $T$-invariant. $T|_{D}\colon D\rightarrow D$ is an open, topologically exact, continuous map that is distance-expanding with respect to the metric $\rho$ with some constants $\eta$, $\lambda$, and $\xi$.
@@
    $\varphi\colon D\rightarrow\mathbb{R}$ and $T|_{D}\colon D\rightarrow\mathbb{R}$ are both computable in the computable metric space $(D,\,\rho,\,\mathcal{S}^{\prime})$.
The Assumption B (ii) through (iv) give the uniqueness of the equilibrium state of $T|_{D}$ for the potential $\varphi$. Together with Theorem 1.1, the Assumption B (ii) through (v) imply that the equilibrium state is computable in the computable metric space $(D,\,\rho,\,\mathcal{S}^{\prime})$. Then the Assumption B (ii) and (iii) allow us to deduce the computability in the computable metric space $(X,\,\rho,\,\mathcal{S})$ from the computability in the computable metric space $(D,\,\rho,\,\mathcal{S}^{\prime})$. Therefore, for the systems satisfying the above assumptions, we prove the following theorem.
@@
Let the septet $(X,\,D,\,\rho,\,\mathcal{S},\,\mathcal{S}^{\prime},\,\varphi,\,T)$ satisfy the Assumption B in Section 1, and $\mu$ be the unique equilibrium state of $T|_{D}$ for the potential $\varphi$. Then $\mu$ is computable.
For a more precise version of Theorem B, see Theorem 1.2 below. As an application, we apply Theorem 1.2 to study the equilibrium states of hyperbolic rational maps on the Riemann sphere associated to the geometric potentials. We say that a rational map is hyperbolic if it is distance-expanding after iterations on the Julia set with respect to the spherical metric. Hyperbolic rational maps are abundant and are expected to be dense among all the rational maps. In the case of quadratic rational maps, see the renowned MLC Conjecture (refer to [Mil06, Appendix G] for reference). The studies on equilibrium states as well as the theory of thermodynamic formalism for hyperbolic rational maps are by now well-developed and established. Given $t\in\mathbb{R}$, it is well-known that a hyperbolic rational map $f$ restricted to the Julia set $\mathcal{J}_{f}$ admits a unique non-atomic equilibrium state $\mu_{t,f}$ which is supported on $\mathcal{J}_{f}$ for the geometric potential $t\varphi_{f}(z)=t\log\bigl{(}f^{\#}(z)\bigr{)}$ (see (5.1) for the definition of $f^{\#}(z)$). With the above conventions, we state the following theorem.
@@
There exists an algorithm which, on input oracles of the coefficients of a hyperbolic rational map $f$ of degree $d\geqslant 2$ and an oracle of $t\in\mathbb{R}$, computes the unique equilibrium state of $f|_{\mathcal{J}_{f}}$ for the potential $t\varphi_{f}(z)$. Moreover, the Hausdorff dimension of the Julia set $\mathcal{J}_{f}$ is computable.
Note that the proof of Theorem C depends on the computability of the hyperbolic Julia sets ([Br04]). Moreover, since the computability of the topological pressure is a consequence of the proof of Theorem 1.1, the pressure function $\mathrm{P}(t)\coloneqq P(f|_{\mathcal{J}_{f}},-t\varphi_{f})$ is computable. Because $\mathrm{P}(t)$ is strictly decreasing with respect to $t\in\mathbb{R}$, as a byproduct, we can compute its unique zero, that is, by Bowen’s formula (see for example, [PU10, Theorem 9.1.6 and Corollary 9.1.7]), the Hausdorff dimension of $\mathcal{J}_{f}$. Non-uniqueness of equilibrium states. Determining which dynamical system admits a unique equilibrium state is a central problem in ergodic theory and a valuable tool for studying the statistical properties of a system (for example, detecting the existence of a phase transition). As indicated by Hofbauer [Ho75], this question was motivated by the studies of intrinsically ergodic dynamical systems, i.e., those having a unique measure of maximal entropy, see for examples, [Bu97, Bu05, Bo74, CT12, Ho79, Ho81, Pa64, Weis70, Weis73]. For general equilibrium states beyond measures of maximal entropy, most results focus on proving the uniqueness, and there are limited examples of systems and potentials with multiple equilibrium states. In the context of a uniformly hyperbolic (or expanding) map, Bowen [Bo75] proved that an equilibrium state does exist and is unique if the potential is Hölder continuous and the map is topologically transitive. In addition, the theory for finite shifts was developed and used to achieve similar results for smooth dynamics. In contrast, Hofbauer [Ho75] constructed a particular class of continuous but non-Hölder potentials $\phi$ for the full shift, each of which admits two equilibrium states. Beyond uniform hyperbolicity, the uniqueness of the equilibrium state was studied by many authors, including Bruin, Keller, Li, Rivera-Letelier, Iommi, Doobs, and Todd [BK98, LRL14, IT15, IT10, DT21] for interval maps; Denker and Urbanski [DU91] for rational maps; the third-named author, Das, Przytycki, Tiozzo, Urbański, and Zdunik [Li18, DPTUZ21] for branched covering maps; Buzzi, Sarig, and Yuri [BS03, Yu03], for countable Markov shifts and for piecewise expanding maps in one and higher dimensions. For local diffeomorphisms with some non-uniform expansion, there are results due to Varandas and Viana [VV10], Pinheiro [Pi11], Ramos and Viana [RV17]. All of these results focus on establishing the existence and uniqueness of the equilibrium state for a potential with low oscillation. On the other hand, much fewer examples that admit multiple equilibrium states have been explored. They include
@@
    Makarov and Smirnov [MS00, MS03], and Rivera-Letelier and Przytycki [PRL11] systematically studied the pressure function for rational maps $f$ with geometric potential $\phi_{t}=-t\log\lvert f^{\#}\rvert$, and constructed examples admitting multiple equilibrium states. For a similar study on (generalized) interval maps, see Rivera-Letelier and Przytycki [PRL14].
@@
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space, and $X$ a recursively compact set. Assume that $T\colon X\rightarrow X$ is a computable map with $h_{\operatorname{top}}(T)=0$ satisfying that there exists no computable point in the non-wandering set $\Omega(T)$. Let $\varphi\colon X\rightarrow\mathbb{R}$ be a computable function satisfying that the pressure $P(T,\varphi)$ is computable. Then there exist at least two equilibrium states of $T$ for $\varphi$.
We refer to (2.1) for the definition of non-wandering sets. For example, it follows immediately from Theorem D and the construction in [GHR10, Subsection 4.1] that given a number $\epsilon>0$, there exists a computable function $T\colon\mathbb{S}^{1}\rightarrow\mathbb{S}^{1}$ from [GHR10, Subsection 4.1] satisfying that for each arc $I$ larger than $\epsilon$ with computable endpoints, there exist at least two equilibrium states of $T$ for the potential $-d_{I}$, where $d_{I}(x)$ denotes the Euclidean distance between $x$ and $I$ along the circle $\mathbb{S}^{1}$ for each $x\in\mathbb{S}^{1}$. More generally, we have the following result.
@@
$\mathcal{H}(\mu)\coloneqq\begin{cases}h_{\mu}(T)&\text{if }\mu\in\mathcal{M}_{T}(X);\\ 0&\text{if }\mu\notin\mathcal{M}_{T}(X)\end{cases}$
is upper-computable on $\mathcal{M}_{T}(X)$, and $\varphi\colon X\rightarrow\mathbb{R}$ is a computable function satisfying that the topological pressure $P(T,\varphi)$ is lower-computable. Then the set consisting of all equilibrium states of $T$ for $\varphi$ is recursively compact. Consequently, if there exists a non-computable equilibrium state of $T$ for $\varphi$, then there exists at least two equilibrium states of $T$ for $\varphi$.
Here $\mathcal{M}_{T}(X)$ is the space of $T$-invariant Borel probability measures equipped with the Wasserstein–Kantorovich metric $W_{\rho}$ (see Section 2.1). Counterexamples. In the following theorem, we provide an example of a computable dynamical system with zero topological entropy (hence, for which equilibrium states exist), whose every equilibrium state for every continuous potential is non-computable.
@@
There exists a computable homeomorphism $T\colon\mathbb{S}^{1}\rightarrow\mathbb{S}^{1}$ of the unit circle $\mathbb{S}^{1}$ such that for each continuous potential $\varphi\colon\mathbb{S}^{1}\rightarrow\mathbb{R}$, the following statements hold:
@@
    There exists at least one equilibrium state of $T$ for $\varphi$.
@@
Recall that the topological entropy reflects a degree of complexity of a dynamical system. We, furthermore, show that there is a dynamical system on the $2$-torus with arbitrarily high topological entropy that admits equilibrium states for each Hölder continuous potential, and whose equilibrium states are all non-computable.
@@
There exists a continuous map $\widehat{T}\colon\mathbb{S}^{1}\times\mathbb{S}^{1}\rightarrow\mathbb{S}^{1}\times\mathbb{S}^{1}$ with non-zero topological entropy such that for each continuous potential $\widehat{\varphi}\colon\mathbb{S}^{1}\times\mathbb{S}^{1}\rightarrow\mathbb{R}$, the following statements hold:
@@
    There exists at least one equilibrium state of $\widehat{T}$ for $\widehat{\varphi}$.

### v3

@@ Abstract.
Equilibrium states are natural dynamical analogues of Gibbs states in thermodynamic formalism. This paper investigates their computability within the framework of Computable Analysis. We show that the unique equilibrium state for a computable, open, topologically exact, distance-expanding map $T\colon X\rightarrow X$ and a computable Hölder continuous potential $\varphi\colon X\rightarrow\mathbb{R}$ is always computable. As an application, we establish the computability of equilibrium states for computable hyperbolic rational maps and their respective geometric potentials. Moreover, we develop a constructive method to exhibit the non-uniqueness of equilibrium states for some dynamical systems. We also present some computable dynamical systems whose equilibrium states are all non-computable.
@@ 1. Introduction
The decidability problem was the third question of Hilbert’s program, which was proposed by Hilbert in Paris in 1900. In the 1930s, Halting Problem given by Turing provided a counterexample to this question. Another significant contribution of Turing was to construct a definite method of computations called Turing machine. Turing machines give a standard model of computation in discrete settings and have become the foundation of modern digital computers.
However, most digital computers are used for computations in continuous settings nowadays. The research on computations with real numbers began with Turing’s work on the original definition of computable real numbers in 1937 ([Tu37]). The work of Banach and Mazur in 1937 ([BM37]) provided definitions of computability for real objects (such as subsets of $\mathbb{R}^{n}$ and functions $f\colon\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$). The “bit model” has been developed, serving as a foundation of the tradition of Computable Analysis ([Gr55, La55, Ko91, Weih00]). Relevant to our paper, in [HR09], Hoyrup and Rojas established a natural computable structure on the set of probability measures, gave a definition of a computable probability measure, and investigated a computable metric space with a computable probability measure.
Computable Analysis on dynamical systems.
In recent decades, there has been dramatic growth in research on the computability and computational complexity of many objects generated by dynamical systems (see e.g. [Ya21] and references therein). The initial value sensitivity and typical instability of many interesting systems imply that what is observed on the computer screen could be completely unrelated to what was meant to be simulated.
This paper is devoted to the study of computability questions for a family of invariant measures called equilibrium states (see Definition 2.5). It expands and continues the line of inquiry initiated by [BBRY11] of the first-named author and Braverman, Rojas, and Yampolsky. Recall that for a rational map f(z)=P(z)/Q(z)f(z)=P(z)/{Q(z)} on the Riemann sphere, where $P(z)$ and $Q(z)$ are mutually prime polynomials, with degree $\deg(f)=\max\{\deg(P),\,\deg(Q)\}\geqslant 2$, the Julia set $\mathcal{J}_{f}$ is defined as the locus of the chaotic dynamics of $f$, namely, the complement of the set where the dynamics of $f$ is stable. Moreover, the weak limit of averages of Dirac measures supported on the preimage points of a repelling periodic point $w$ is an $f$-invariant measure $\lambda$ (independent of the choice of $w$), called Brolin–Lyubich measure.
In [BBRY11], a uniform machine was designed to compute the Brolin–Lyubich measure for any given rational map. Together with the existence of polynomials whose coefficients are all computable but whose Julia sets are non-computable from the groundbreaking works of Braverman and Yampolsky ([BY06, BY09]) dating back to a question of Milnor [BY06, Section 1], the computability of Brolin–Lyubich measures leads to a conflict: heuristically speaking, a measure contains more information than its support, but in Computable Analysis, there exists a computable invariant probability measure whose support is, however, non-computable. Such a conflict can be reconciled by considering these two results as the computability properties of the same physical objects from the geometric and statistical perspectives, respectively. In this era of artificial intelligence and data science, the latter undoubtedly deserves closer investigation.
Regarding the algorithmic aspects of general invariant measures, in a fundamental paper [GHR11], Galatolo, Hoyrup, and Rojas established that for each computable map $T\colon X\rightarrow X$ on a recursively compact set $X$, all isolated invariant measures are computable. As a counterexample, a computable dynamical system with no computable invariant measures is constructed in [GHR11, Subsection 4.1] as well. In Sections 4 and 7, we extend these results.
Thermodynamic formalism is a powerful method for investigating invariant measures with prescribed local behavior under iterations of the dynamical system. This theory, inspired by statistical mechanics, was created by Ruelle, Sinai, and others in the seventies ([Dob68, Si72, Bo75, Wa82]). Since then, thermodynamic formalism has been applied in many classical contexts (see e.g. [Bo75, Ru89, Pr90, KH95, Zi96, MauU03, BS03, Ol03, Yu03, PU10, MayU10]) and has remained at the forefront of research in dynamical systems.
Among many other applications, thermodynamic formalism has played a central role in the study of the statistical properties of dynamical systems since its very early days.
The key objects of investigation in thermodynamic formalism are invariant measures called equilibrium states. More precisely, for a continuous map $T\colon X\rightarrow X$ of a compact metric space $(X,\rho)$, a $T$-invariant Borel probability measure $\mu$ on $X$, and a real-valued continuous function $\varphi\colon X\rightarrow\mathbb{R}$, called potential, one can define the measure-theoretic pressure and the topological pressure to describe the degree of complexity of the given dynamical system from different perspectives. The Variational Principle implies that the supremum over all $T$-invariant Borel probability measures on $X$ of the measure-theoretic pressure is equal to the topological pressure. We say that a Borel probability measure $\mu$ is an equilibrium state for $T$ and $\varphi$ if $\mu$ maximizes the measure-theoretic pressure. In particular, if $\varphi$ is a constant function, then an equilibrium state reduces to a measure of maximal entropy (see Subsection 2.2). In particular, the Brolin–Lyubich measure of a rational map $f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ is the measure of maximal entropy of $f$.
In recent years, significant progress has been made in understanding the computability of dynamical invariants, such as topological entropy and pressure (see e.g. [Sp07, Sp08, BDWY22]). In the proof of Theorem A, we use a novel approach to establish the computability of the topological pressure and the computability of the Jacobian function in a different setting.
Statement of main results.
Our main contributions are threefold: First, we establish computability results for equilibrium states in uniformly expanding systems (see Theorems A and B) and apply them to demonstrate the computability of equilibrium states for some complex dynamical systems (see Theorem C). Second, we introduce a novel mechanism to demonstrate non-uniqueness of equilibrium states (see Theorems D and E). Third, we construct computable systems with non-computable equilibrium states (see Theorems F and G).
Computability of equilibrium states.
Our first result concerns the computability of equilibrium states. Here is a list of assumptions used in this context.
@@
    $\varphi\colon X\rightarrow\mathbb{R}$ is a Hölder continuous function with an exponent $v_{0}\in(0,1]$.
@@
    $T\colon X\rightarrow X$ is an open, topologically exact, continuous, and distance-expanding map with respect to the metric $\rho$ with constants $\eta>0$, $\lambda>1$, and $\xi>0$ (in the sense of Definition 2.1).
@@
    $\varphi$ and $T$ are both computable (in the sense of Definition 3.9).
By classical results (see e.g. Proposition 2.6 and [PU10, Corollary 5.2.14]), the Assumption A (i) through (iii) imply that there exists a unique equilibrium state for $T$ and $\varphi$. The Assumption A (iv) allows us to input the information of the system in the computer.
@@
Let the quintet $(X,\,\rho,\,\mathcal{S},\,\varphi,\,T)$ satisfy the Assumption A in Section 1, and $\mu$ be the unique equilibrium state for the map $T$ and the potential $\varphi$. Then $\mu$ is computable.
For a more detailed version of Theorem A, see Theorem 1.1 below. The Assumption A requires that the behavior of the dynamical system has nice properties in the whole space. However, the dynamical behavior may be good enough in an invariant subspace rather than in the whole space. Hence, we prove Theorem B, a “subspace” version of Theorem A. We give a list of assumptions that describe the systems considered in Theorem B as follows.
@@
    $(D,\,\rho^{\prime},\,\mathcal{S}^{\prime})$ is a computable metric space, where $D$ is a subset of $X$, $\rho^{\prime}$ is the restriction of the metric $\rho$ on $D$, and $\mathcal{S}^{\prime}$ is a uniformly computable sequence of points in the computable metric space $(X,\,\rho,\,\mathcal{S})$. The subset $D$ is a recursively compact set in the computable metric space $(D,\,\rho^{\prime},\,\mathcal{S}^{\prime})$.
@@
    $\varphi\colon D\rightarrow\mathbb{R}$ is a Hölder continuous function with an exponent $v_{0}\in(0,1]$.
@@
    $T\colon X\rightarrow X$ is a map such that $D$ is completely $T$-invariant. $T|_{D}\colon D\rightarrow D$ is an open, topologically exact, continuous map that is distance-expanding with respect to the metric $\rho^{\prime}$ with constants $\eta>0$, $\lambda>1$, and $\xi>0$.
@@
    $\varphi\colon D\rightarrow\mathbb{R}$ and $T|_{D}\colon D\rightarrow\mathbb{R}$ are both computable in the computable metric space $(D,\,\rho^{\prime},\,\mathcal{S}^{\prime})$.
The Assumption B (ii) through (iv) give the uniqueness of the equilibrium state for the restriction $T|_{D}$ and the potential $\varphi$. Together with Theorem 1.1, the Assumption B (ii) through (v) imply that the equilibrium state is computable in the computable metric space $(D,\,\rho^{\prime},\,\mathcal{S}^{\prime})$. Then the Assumption B (ii) and (iii) allow us to deduce the computability in the computable metric space $(X,\,\rho,\,\mathcal{S})$ from the computability in the computable metric space $(D,\,\rho^{\prime},\,\mathcal{S}^{\prime})$. Therefore, for the systems that satisfy the above assumptions, we prove the following theorem.
@@
Let the septet $(X,\,D,\,\rho,\,\mathcal{S},\,\mathcal{S}^{\prime},\,\varphi,\,T)$ satisfy the Assumption B in Section 1, and $\mu$ be the unique equilibrium state for the restriction $T|_{D}$ and the potential $\varphi$. Then $\mu$ is computable.
For a more detailed version of Theorem B, see Theorem 1.2 below. We apply Theorem B (or Theorem 1.2) to establish Theorem C, i.e., the computability of the equilibrium states for the restriction of hyperbolic rational maps on their corresponding Julia sets and their corresponding geometric potentials. Indeed, there are many other dynamical systems whose dynamical behavior may be merely good enough in an invariant subspace.
We say that a rational map is hyperbolic if its postcritical orbit closure is disjoint from its Julia set. Hyperbolic rational maps are abundant and are expected to be dense among all rational maps. In the case of quadratic rational maps, see the renowned MLC Conjecture; namely, the Mandelbrot set is locally connected (refer to [Mil06, Appendix G] for reference). The studies on equilibrium states as well as the theory of thermodynamic formalism for hyperbolic rational maps are by now well-developed and established. Given $t\in\mathbb{R}$, it is well-known that a hyperbolic rational map $f$ restricted to the Julia set $\mathcal{J}_{f}$ admits a unique non-atomic equilibrium state $\mu_{t,f}$ which is supported on $\mathcal{J}_{f}$ for the geometric potential $t\varphi_{f}(z)=t\log\bigl{(}f^{\#}(z)\bigr{)}$ (see (5.1) for the definition of the spherical derivative $f^{\#}(z)$).
With the above conventions, we state the following theorem.
@@
There exists an algorithm which, on input oracles of the coefficients of a hyperbolic rational map $f$ of degree $d\geqslant 2$ and an oracle of $t\in\mathbb{R}$, computes the unique equilibrium state for $f|_{\mathcal{J}_{f}}$ and $t\varphi_{f}(z)$.
Note that the proof of Theorem C depends on the computability of the hyperbolic Julia sets ([Br04]). Moreover, since the computability of the topological pressure is a consequence of the proof of Theorem 1.1, the pressure function $\mathrm{P}(t)\coloneqq P\bigl{(}f|_{\mathcal{J}_{f}},-t\varphi_{f}\bigr{)}$ is computable. Because $\mathrm{P}(t)$ is strictly decreasing with respect to $t\in\mathbb{R}$, as a byproduct, we can compute its unique zero, that is, by Bowen’s formula (see e.g. [PU10, Theorem 9.1.6 and Corollary 9.1.7]), the Hausdorff dimension of $\mathcal{J}_{f}$. For related research on the computation of Hausdorff dimensions of invariant sets, see also [JP02, JP04] and the references therein.
Notably, our approach can be further developed to establish the computability of equilibrium states for more general uniformly hyperbolic systems or even partially hyperbolic systems. Due to space limitations, we focus on the current setup in this paper and postpone further investigation to future works.
Non-uniqueness of equilibrium states.
Determining which dynamical system admits a unique equilibrium state is a central problem in ergodic theory and a valuable tool for studying the statistical properties of a system (for example, detecting the existence of a phase transition).
As indicated by Hofbauer [Ho77], this question was motivated by the studies of intrinsically ergodic dynamical systems, i.e., those having a unique measure of maximal entropy, see e.g. [Bu97, Bu05, Bo74, CT12, Ho79, Ho81, Pa64, Weis70, Weis73].
For general equilibrium states beyond measures of maximal entropy, most results focus on proving the uniqueness, and there are limited examples of systems and potentials with multiple equilibrium states. In the context of a uniformly hyperbolic (or expanding) map, Bowen [Bo75] proved that an equilibrium state exists and is unique if the potential is Hölder continuous and the map is topologically transitive. In addition, the theory for finite shifts was developed and used to achieve similar results for smooth dynamics. In contrast, Hofbauer [Ho77] constructed a particular class of continuous but non-Hölder potentials for the full shift, each of which admits two equilibrium states.
Beyond uniform hyperbolicity, the uniqueness of the equilibrium state was studied by many authors, including Bruin, Keller, Li, Rivera-Letelier, Iommi, Doobs, and Todd [BK98, LRL14, IJT15, IT10, DT23] for interval maps; Denker and Urbański [DU91] for rational maps; the third-named author, Das, Przytycki, Tiozzo, Urbański, and Zdunik [Li18, DPTUZ21] for branched covering maps; Buzzi, Sarig, and Yuri [BS03, Yu03] for countable Markov shifts and for piecewise expanding maps in one and higher dimensions. For local diffeomorphisms with some non-uniform expansion, there are results due to Varandas and Viana [VV10], Pinheiro [Pi11], Ramos and Viana [RV17]. All of these results focus on establishing the existence and uniqueness of the equilibrium state for a potential with low oscillation.
On the other hand, much fewer examples that admit multiple equilibrium states have been explored. They include
@@
    Makarov and Smirnov [MS00, MS03], and Rivera-Letelier and Przytycki [PRL11] systematically studied the pressure function for rational maps $f$ with geometric potential $\phi_{t}=-t\log\bigl{(}f^{\#}\bigr{)}$, and constructed examples admitting multiple equilibrium states. For a similar study on (generalized) interval maps, see Rivera-Letelier and Przytycki [PRL14].
@@
$\mathcal{H}(\mu)\coloneqq\begin{cases}h_{\mu}(T)&\text{if }\mu\in\mathcal{M}(X,T);\\ 0&\text{if }\mu\notin\mathcal{M}(X,T)\end{cases}$
is upper-computable on the space $\mathcal{M}(X,T)$ of $T$-invariant Borel probability measures, and $\varphi\colon X\rightarrow\mathbb{R}$ is a computable function satisfying that $P(T,\varphi)$ is lower-computable. Then the set $\mathcal{E}(T,\varphi)$ of equilibrium states for $T$ and $\varphi$ is recursively compact. Consequently, if there exists a non-computable equilibrium state for $T$ and $\varphi$, then there exist at least two equilibrium states for $T$ and $\varphi$.
As a corollary, we obtain the following result under easier-to-formulate conditions. We refer to (2.2) for the definition of non-wandering sets.
@@
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space, and $X$ a recursively compact set. Assume that $T\colon X\rightarrow X$ is a computable map with $h_{\operatorname{top}}(T)=0$ satisfying that there exists no computable point in the non-wandering set $\Omega(T)$. Let $\varphi\colon X\rightarrow\mathbb{R}$ be a computable function satisfying that the pressure $P(T,\varphi)$ is computable. Then there exist at least two equilibrium states for $T$ and $\varphi$.
For example, the computable map $T\colon\mathbb{S}^{1}\rightarrow\mathbb{S}^{1}$ constructed in [GHR11, Subsection 4.1] satisfies that $h_{\operatorname{top}}(T)=0$ and $\Omega(T)$ contains no computable point. Then for each computable function $\varphi\colon X\rightarrow\mathbb{R}$ satisfying that $P(T,\varphi)$ is computable, there exist at least two equilibrium states for the map $T$ and the potential $\varphi$.
Counterexamples.
In [GHR11], a computable system with no computable invariant measure is constructed. On the basis of this, we establish the following statement.
@@
There exists a computable homeomorphism $T\colon\mathbb{S}^{1}\rightarrow\mathbb{S}^{1}$ of the unit circle $\mathbb{S}^{1}$ (such as the map $T$ constructed in [GHR11, Subsection 4.1]) with the following property: for each continuous potential $\varphi\colon\mathbb{S}^{1}\rightarrow\mathbb{R}$,
@@
    there exists at least one equilibrium state for $T$ and $\varphi$,
@@
Recall that the topological entropy reflects a degree of complexity of a dynamical system. We, furthermore, show that there is a dynamical system on the $2$-torus with arbitrarily high topological entropy that admits equilibrium states for each continuous potential and whose equilibrium states are all non-computable. More precisely, define $\widehat{T}\colon\mathbb{S}^{1}\times\mathbb{S}^{1}\rightarrow\mathbb{S}^{1}\times\mathbb{S}^{1}$ by $\widehat{T}(x,y)=(T(x),T_{d}(y))$ for each $x,\,y\in\mathbb{S}^{1}$, where $T_{d}(x)\coloneqq dx$ for $d\in\mathbb{N}$ and $x\in\mathbb{S}^{1}.$
@@
The map $\widehat{T}\colon\mathbb{S}^{1}\times\mathbb{S}^{1}\rightarrow\mathbb{S}^{1}\times\mathbb{S}^{1}$ constructed above has non-zero topological entropy and satisfies that for each continuous potential $\widehat{\varphi}\colon\mathbb{S}^{1}\times\mathbb{S}^{1}\rightarrow\mathbb{R}$, the following statements are true:
@@
    There exists at least one equilibrium state for $\widehat{T}$ and $\widehat{\varphi}$.